# Mitigating the Human-Robot Domain Discrepancy in Visual Pre-training for Robotic Manipulation

### v1

@@ 4.1 Experimental Setups
The experimental validation includes two stages. First, given existing human-data pre-trained models, we adapt the models for mitigating the human-robot domain discrepancy using the proposed HR-Align method. Second, the pre-trained models after adaptation are used as frozen visual backbones for learning manipulation policies in downstream tasks.
@@
Pre-trained Model Adaptation. To comprehensively evaluate the efficacy of the proposed adaptation method, two existing models pre-trained on different human data for robotic manipulation are selected for adaptation. One is the R3M model [1] pre-trained on Ego4D data [14] with vision-language contrastive pre-training. The other one comes from data4robotics [11], which is pre-trained on Kinetics [70] with MoCo pre-training [71] (We refer to it as D4R model). For the paired human-robot video data used in our HR-Align method, we instantiate it using the existing RH20T dataset [72]. We select a subset from the RH20T dataset, which includes a total of $56$k human-robot video pairs. We will release the dataset profile used in this work. During adaptation, we randomly sample $5$ frames from each video (i.e., $T=5$). Following R3M [1], we use the frozen DistilBert model [69] to process the task descriptions. The Adam optimizer [73] with the learning rate of $1\times e^{-4}$ and batch size of $200$ is used during training. The temperature factor $\tau$ is set to $0.1$. For the adapter module, we use group convolution to instantiate it. The adapter module can be inserted into any position of the visual backbone network. Without specific statements, we insert the adapter module after the last layer of the backbone network. $4$ NVIDIA A6000 GPUs are used in experiments.
@@
Downstream Policy Learning. To learn manipulation policies for different downstream tasks, the pre-trained model after adaptation is used to extract the state representations of the robot’s visual observations. We evaluate a total of 25 robotic manipulation tasks from three different environments, including 7 single-tasks and 18 language-conditioned multi-tasks.
@@
(i) For the single-task setting, following previous works [1; 10], two dexterous manipulation tasks from Adroit [35], and five tasks with a Sawyer robot from Metaworld [34] are evaluated. To learn the motor-control policy for each task, we train the behavior cloning networks with 3 different seeds, using one viewpoint and the largest demo dataset size used in [1]. We report the average success rate over 3 seeds on each task. All other configurations (e.g., network designs, training steps, and evaluation interval) are the same as those in R3M [1].
@@
(ii) For the multi-task setting, we evaluate our adapted pre-trained models on 18 multi-tasks from RLBench [28]. Existing pre-training works [1; 9; 10; 11; 62] for robotic manipulation only evaluate models under single-task setting, where the data scale, task diversity, and task complexity are limited. By additionally evaluating our method on the RLBench, which is challenging for its multi-task setting with significant task diversity and difficulty, the generalizability of our adapted models can be more comprehensively demonstrated. Similar to the data processing pipeline in RVT [7], we discard all the pooling layers in our pre-trained models so that the spatial dimension of the feature map is the same as that of the original image. Instead of using 8 self-attention layers like RVT [7], we only use one single self-attention layer to fuse the frozen image features and language features, and train the network for five epochs and test the network three times. Following RVT [7], we use a total of 1800 demonstrations and report the average success rate and standard deviation. All other experimental configurations are the same as RVT [7]. Please see details of our downstream policy designs on RLBench in Section S2 of the Appendix.
@@ 4.2 Does the HR-Align method enable better generalization across different tasks?
To achieve better generalization on learning downstream manipulation policies, we adapt two human-data pre-trained models, namely, D4R [11] and R3M [1]. We denote the pre-trained models after adaptation as D4R-Align and R3M-Align, respectively.
Table 1: Success rate of two tasks in Adroit. Two pre-trained models, i.e., D4R and R3M, are adapted by our method. The adapted D4R-Align and R3M-Align models achieve better performance on both tasks compared to their unadapted counterparts.
@@
Single-task setting. As shown in Table 1, we evaluate our adapted models on two tasks in the Adroit environment. For the D4R model, our adapted version, namely the D4R-Align model, achieves improvements on both tasks, with an average success rate increased by $2.0\%$ compared to the unadapted D4R model. For the R3M model that was pre-trained on different human data with different pre-trained methodologies, our R3M-Align model also improves the success rate of both tasks, especially for the relocate task (one of the most difficult tasks in Adroit). On average, our R3M-Align model has a significant $7.3\%$ improvement in success rate over the unadapted R3M model.
In addition to the Adroit environment, we also evaluate our method on five tasks from the Metaworld environment. As shown in Table 2, compared with the pre-trained D4R model, our adapted D4R-Align model achieves significant improvements on two out of five tasks. And the average success rate of all five tasks is improved by $2.2\%$ in this challenging environment. For the pre-trained R3M model, our adapted R3M-Align model achieves significant improvements on three tasks and comparable performance on the other two tasks. On average, the adapt R3M-Align model has $2.3\%$ improvement in the average success rate of five tasks.
Table 2: Success rate of five tasks in Metaworld. Compared to the two unadapted pre-trained models (i.e., D4R and R3M), our adapted D4R-Align and R3M-Align models achieve satisfactory improvements in the averaged success rate of the five tasks.
Models assembly bin-pick button-press drawer-open hammer Averaged
D4R 65.3 78.7 60.7 100.0 88.0 78.5
D4R-Align 80.7 75.3 56.7 100.0 90.7 80.7 (+2.2)
R3M 78.7 53.3 60.0 100.0 94.0 77.2
R3M-Align 82.0 58.0 60.7 100.0 96.7 79.5 (+2.3)
Multi-task setting. Different from previous works that only evaluate models under the above single-task setting, to more comprehensively evaluate the effectiveness of our method, we evaluate the adapted pre-trained models on $18$ language-conditioned tasks from the RLBench benchmark. This multi-task setting is challenging, as it requires learning a single policy to complete various instructions in all $18$ manipulation tasks. For each task, we report the average success rate and standard deviation over three tests. The results are shown in the Table 3. Compared to the unadapted D4R model, our adapted D4R-Align model achieves significantly superior performance on $10$ tasks, with $4.6\%$ improvement in the averaged success rate of all $18$ tasks. Similarly, our adapted R3M-Align model also significantly improves the success rate of $10$ tasks compared to the unadapted R3M model. More importantly, on this challenging benchmark, our adapted R3M-Align model achieves $8.9\%$ improvement in the averaged success rate of all $18$ tasks, demonstrating the necessity of mitigating the human-robot domain discrepancy in motor-control pre-training.
Table 3: Success rate of 18 tasks in RLBench. Compared to the two unadapted D4R and R3M models, our adapted D4R-Align and R3M-Align models significantly improve the success rate of most tasks. The average improvement of each model on all 18 tasks is also significant.
@@
D4R-Align $\textbf{32.0}_{\pm 4.0}$ $97.3_{\pm 2.3}$ $96.0_{\pm 0.0}$ $\textbf{70.7}_{\pm 2.3}$ $81.3_{\pm 2.3}$ $\textbf{46.7}_{\pm 4.6}$ $22.7_{\pm 2.3}$ $\textbf{97.3}_{\pm 2.3}$ $\textbf{100.0}_{\pm 0.0}$ $\textbf{72.0}_{\pm 6.9}$
@@
R3M-Align $\textbf{70.7}_{\pm 6.1}$ $\textbf{97.3}_{\pm 2.3}$ $\textbf{80.0}_{\pm 4.0}$ $\textbf{80.0}_{\pm 4.0}$ $\textbf{80.0}_{\pm 8.0}$ $\textbf{33.3}_{\pm 9.2}$ $13.3_{\pm 2.3}$ $78.7_{\pm 4.6}$ $\textbf{100.0}_{\pm 0.0}$ $\textbf{80.0}_{\pm 4.0}$
@@
D4R-Align $\textbf{16.0}_{\pm 4.0}$ $0.0_{\pm 0.0}$ $\textbf{84.0}_{\pm 4.0}$ $\textbf{57.3}_{\pm 8.3}$ $\textbf{86.7}_{\pm 4.6}$ $24.0_{\pm 6.9}$ $92.0_{\pm 4.0}$ $\textbf{1.3}_{\pm 2.3}$ 59.9 (+4.6)
@@
R3M-Align $13.3_{\pm 2.3}$ $0.0_{\pm 0.0}$ $\textbf{88.0}_{\pm 4.0}$ $\textbf{41.3}_{\pm 4.6}$ $\textbf{98.7}_{\pm 2.3}$ $16.0_{\pm 6.9}$ $93.3_{\pm 2.3}$ $\textbf{1.3}_{\pm 2.3}$ 59.2 (+8.9)
@@ 4.4 How much does each component contribute to the HR-Align method?
Driven by the idea of human-robot semantic alignment, another contribution of this work is leveraging the parameter-efficient design (i.e., adapter) to adapt the representation of pre-trained models. Following existing practice [22; 24; 25], for a visual backbone consisting of multiple blocks (e.g., the residual blocks in resnet [74]), we selectively add the adapter modules at three positions in the backbone network. Taking the pre-trained R3M model as an example. We can insert our adapter modules before the first block, in the middle of all blocks, and after the last block. We call these three adapted models R3M-Align-E, R3M-Align-M, and R3M-Align-L, respectively. Similarly, we also add the adapter modules at all three locations of the R3M model, which we denote as R3M-Align-E.M.L.
@@
In this work, we also use the task description features as queries to better capture manipulation-oriented semantics from video features. As shown at the bottom of Table 5, removing this operation from the adapted model, i.e., R3M-Align-L w/o lang., will result in performance degradation. This demonstrates that language-guided feature enhancement is effective in promoting human-robot semantic alignment.
Table 5: Success rates of two tasks in Adroit. Middle: ablation on inserting adapter modules at different positions in the R3M model. Bottom: ablation on the language-guided feature enhancement.
@@
4.5 Is the adapted representation more effective for downstream tasks?
To investigate the effectiveness of our proposed adaptation method, we employ t-SNE [75] to depict the downstream feature distributions of the existing pre-trained R3M model and our adapted R3M-Align model, respectively. We randomly select 5 tasks from the RLBench dataset, where 500 samples are randomly selected for each task. The distributions of feature representation of the R3M model and R3M-Align model are shown in Figure 3(a) and Figure 3(b), respectively. In the figure, different colors represent different tasks, and each point represents a sample.
In the visualization, we find that the feature distribution of the existing R3M model is significantly more dispersed than that of our adapted R3M-Align model. This phenomenon demonstrates that our adaptation method can effectively learn more discriminative representation for downstream tasks by addressing the human-robot domain discrepancy in pre-training.
5 Conclusion
The existing learning paradigm of visual pre-training on human data for robotic manipulation encounters the human-robot domain discrepancy. Our work takes a preliminary attempt to solve this challenging problem. In this work, we contribute a new adaptation paradigm by leveraging existing semantic-aligned human-robot video data and proposing an efficient semantic alignment method. In this way, the existing human-data pre-trained models can be efficiently and explicitly adapted to the robot domain, without the need to be tailored for each downstream robotic environment. Experiments on 25 robotic manipulation tasks across three environments and different pre-trained models demonstrate the efficacy of our proposed method.
Appendix
Mitigating the Human-Robot Domain Discrepancy
in Visual Pre-training for Robotic Manipulation
Appendix S1 Broader Impacts Statements
There are no ethical issues involved in this paper. This paper presents work whose goal is to advance the field of robotic learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix S2 Downstream Policy Learning on RLBench
Existing visual pre-training works often assess the learned representation in downstream environments using a single-task setting. However, our work diverges from this standard by also evaluating the adapted pre-trained models on the large-scale RLBench benchmark [28], where a single language-conditioned policy is learned to complete various tasks.
On RLBench, existing works [6; 7; 12; 32; 33] usually develop sophisticated models for modeling the semantics of robot’s multi-view observations and their correlations with the language commands. For example, RVT [7] utilizes four attention layers as the visual encoder to model intra-image relations and four more attention layers to model image-language correlations. In this work, we adopt the same design as RVT, but replace the visual encoder (i.e., four intra-image attention layers) with either an existing pre-trained model or our adapted one. Furthermore, we employ just one attention layer to fuse the extracted image features and language features. As RVT predicts the end-effector’s actions based on the features without down-sampling the spatial dimension, we discard all spatial down-sampling operations (e.g., max-pooling) in the pre-trained models and our adapted models. Please note that to validate the effectiveness of our adaptation method, we freeze the visual representation of the pre-trained model or our adapted model while learning the downstream policy on RLBench.

### v3

@@ 4.1 Experimental Setups
The experimental validation includes two stages. First, given existing human-data pre-trained models, we adapt the models for mitigating the human-robot domain discrepancy using the proposed HR-Align method. Second, the pre-trained models after adaptation are used as frozen visual backbones for learning manipulation policies on downstream tasks.
@@
Pre-trained Model Adaptation. To comprehensively evaluate the efficacy of the proposed adaptation method, we adapt two human-data pre-trained models that are widely used in the robotic manipulation community. One is the R3M model [50] pre-trained on Ego4D [21] with vision-language contrastive pre-training. The other one comes from data4robotics [13], which is pre-trained on Kinetics [7] with MoCo pre-training [27] (We refer to it as D4R model). For the paired human-robot video data, we instantiate it using the existing RH20T dataset [16]. We select a subset from the RH20T dataset, which includes a total of $56$k human-robot video pairs. During adaptation, we randomly sample $5$ frames from each video (i.e., $T=5$). Following R3M [50], we use the frozen DistilBert model [54] to process the task descriptions. The Adam optimizer [37] with the learning rate of $1\times e^{-4}$ and batch size of $200$ is used during training. The temperature factor $\tau$ is set to $0.1$. Without specific statements, we insert the adapter module after the last layer of the backbone network. The adaptation process takes about $8$k training steps with $4$ NVIDIA A6000 GPUs, requiring a small extra cost compared to human-data pre-training.
@@
Downstream Policy Learning. To learn manipulation policies for different downstream tasks, the pre-trained model after adaptation is used to extract the state representations of the robot’s visual observations. Then, action prediction models are built upon these frozen visual features to predict the next position of the robot’s end-effector.
Figure 3: Left: our real-world experimental workspace. Top right: illustrations of five real-world tasks. Bottom right: experimental results on five tasks, where the pre-trained D4R and R3M, and our adapted D4R-Align, R3M-Align models are evaluated.
(i) Simulated tasks. For simulation environments, we evaluate a total of 20 robotic manipulation tasks across two different simulated benchmarks, including 2 single-tasks and 18 language-conditioned multi-tasks.
@@
For the single-task setting, following previous works [50, 29], two dexterous manipulation tasks from Adroit [53] are evaluated. To learn the motor-control policy for each task, we train the behavior cloning networks with 3 different seeds, using the top-view inputs and the largest demo dataset size used in [50]. We report the average success rate of over 3 seeds for each task. All other configurations (e.g., network designs, training steps, and evaluation interval) are the same as those in R3M [50].
@@
For the multi-task setting, we evaluate our adapted pre-trained models on 18 multi-tasks from RLBench [32]. Existing pre-training works [50, 68, 29, 13, 33] for robotic manipulation only evaluate models under single-task setting, where the data scale, task diversity, and task complexity are limited. By additionally evaluating our method on the RLBench, which is challenging for its multi-task setting with significant task diversity and difficulty, the generalizability of our adapted models can be more comprehensively demonstrated. Instead of using 8 self-attention layers like RVT [19], we only use one single self-attention layer to fuse the frozen image features and language features, and train the network for five epochs and test the network three times. Following RVT [19], we use a total of 1800 demonstrations for training and report the average success rate and standard deviation for testing. All other experimental configurations are the same as RVT [19]. Please see details of our downstream policy designs on RLBench in the appendix.
(ii) Real-world tasks. For the real-world environment, we use an xArm7 robot arm, an Inspire gripper, and a third-view Orbbec Femto Bolt camera. Five tasks, i.e., put fruit in plate, stack cups, put tennis in mug, hang mug, and put item in box are designed. Figure 3 illustrates our experimental setups. For each task, we collect 40 demonstrations for training. During evaluation, we test 20 episodes for each task and report the average success rate. For simplicity, we train the real-world policy under a single task setting, where the pre-trained models or our adapted models are used as visual feature extractors, with an ACT [76] framework to predict the following key-actions (likes RVT [19]) of the end-effector. For more details about the setups, data collection, and model designs, please refer to the appendix.
@@ 4.2 Does the HR-Align method enable better generalization across different tasks?
To demonstrate the effectiveness of our adaptation method on downstream tasks, we adapt two human-data pre-trained models that are widely used in the community, namely, D4R [13] and R3M [50]. We denote the pre-trained models after adaptation as D4R-Align and R3M-Align, respectively.
Simulated single-task setting. As shown in Table 1, we evaluate our adapted models on two tasks in the Adroit environment. For the D4R model, our adapted version, namely the D4R-Align model, achieves improvements on both tasks, with an average success rate increased by $2.0\%$ compared to the unadapted D4R model. For the R3M model that was pre-trained on different human data with different pre-trained methodologies, our R3M-Align model also improves the success rate of both tasks, especially for the relocate task (one of the most difficult tasks in Adroit). On average, our R3M-Align model has a significant $7.3\%$ improvement in success rate over the unadapted R3M model.
@@
Table 1: Success rate of tasks in Adroit. Two pre-trained models, i.e., D4R and R3M, are adapted by our method. The adapted D4R-Align and R3M-Align models achieve better performance on the tasks compared to their unadapted counterparts.
@@
D4R-Align $32.0_{\pm 4.0}$ $97.3_{\pm 2.3}$ $96.0_{\pm 0.0}$ $70.7_{\pm 2.3}$ $81.3_{\pm 2.3}$ $46.7_{\pm 4.6}$ $22.7_{\pm 2.3}$ $97.3_{\pm 2.3}$ $100.0_{\pm 0.0}$ $72.0_{\pm 6.9}$
@@
R3M-Align $70.7_{\pm 6.1}$ $97.3_{\pm 2.3}$ $80.0_{\pm 4.0}$ $80.0_{\pm 4.0}$ $80.0_{\pm 8.0}$ $33.3_{\pm 9.2}$ $13.3_{\pm 2.3}$ $78.7_{\pm 4.6}$ $100.0_{\pm 0.0}$ $80.0_{\pm 4.0}$
@@
D4R-Align $16.0_{\pm 4.0}$ $0.0_{\pm 0.0}$ $84.0_{\pm 4.0}$ $57.3_{\pm 8.3}$ $86.7_{\pm 4.6}$ $24.0_{\pm 6.9}$ $92.0_{\pm 4.0}$ $1.3_{\pm 2.3}$ $59.9~{}{\color[rgb]{1.0,0.0,0.5}(+4.6)}$
@@
R3M-Align $13.3_{\pm 2.3}$ $0.0_{\pm 0.0}$ $88.0_{\pm 4.0}$ $41.3_{\pm 4.6}$ $98.7_{\pm 2.3}$ $16.0_{\pm 6.9}$ $93.3_{\pm 2.3}$ $1.3_{\pm 2.3}$ $59.2~{}{\color[rgb]{1.0,0.0,0.5}(+8.9)}$
Table 2: Success rate of 18 tasks in RLBench. Compared to the unadapted D4R and R3M models, our adapted D4R-Align and R3M-Align models significantly improve the success rate of most tasks. The average improvement of each model on all 18 tasks is also significant.
Simulated multi-task setting. Unlike previous works that evaluate models solely under the single-task setting, we conduct a more comprehensive evaluation of our method by evaluating the adapted pre-trained models on 18 language-conditioned tasks from the RLBench benchmark. This multi-task setting is challenging, as it requires learning a single policy to complete various instructions in all $18$ manipulation tasks. For each task, we report the average success rate and standard deviation over three tests. The results are shown in the Table 2. Compared to the unadapted D4R model, our adapted D4R-Align model achieves significantly superior performance on $10$ tasks, with $4.6\%$ improvement in the averaged success rate of all $18$ tasks. Similarly, our adapted R3M-Align model also significantly improves the success rate of $10$ tasks compared to the unadapted R3M model. More importantly, on this challenging benchmark, our adapted R3M-Align model achieves $8.9\%$ improvement in the averaged success rate of all $18$ tasks, demonstrating the necessity of mitigating the human-robot domain discrepancy in motor-control pre-training.
Real-world task setting. In the bottom right of Figure 3, we present the performance of the pre-trained D4R and R3M models, as well as our adapted D4R-Align and R3M-Align models, across the five tasks. For each task, the results indicate that our adapted models achieve significant improvements over their unadapted versions. On average, across all five tasks, the adapted D4R-Align model shows a $13\%$ improvement in success rate compared to the pre-trained D4R model, while the adapted R3M-Align model similarly demonstrates an $11\%$ improvement over the pre-trained R3M model. These results underscore the effectiveness of our adaptation method in improving existing visual pre-trained models in real-world scenarios.
@@ 4.4 How much does each component contribute to the HR-Align method?
Driven by the idea of human-robot semantic alignment, another contribution of this work is leveraging the parameter-efficient design (i.e., adapter) to adapt the representation of pre-trained models. Following existing practice [58, 73, 51], for a visual backbone consisting of multiple blocks (e.g., the residual blocks in resnet [25]), we selectively add the adapter modules at three positions in the backbone network. Taking the pre-trained R3M model as an example. We can insert our adapter modules before the first block, in the middle of all blocks, and after the last block. We call these three adapted models R3M-Align-E, R3M-Align-M, and R3M-Align-L, respectively. Similarly, we also add the adapter modules at all three locations of the R3M model, which we denote as R3M-Align-E.M.L.
@@
Table 4: Success rates of two tasks in Adroit. Ablation studies on inserting adapter modules at different positions in the R3M model.
4.5 Is the adapted representation more effective on downstream tasks?
To investigate the effectiveness of our proposed adaptation method, we employ t-SNE [62] to depict the downstream feature distributions of the existing pre-trained R3M model and our adapted R3M-Align model, respectively. We randomly select 5 tasks from the RLBench dataset, where 500 samples are randomly selected for each task. The distributions of feature representation of the R3M model and R3M-Align model are shown in Figure 4 (a) and Figure 4 (b), respectively. In the figure, different colors represent different tasks, and each point represents a sample. In the visualization, we find that the feature distribution of the existing R3M model is significantly more dispersed than that of our adapted R3M-Align model. This phenomenon demonstrates that our adaptation method can effectively learn more discriminative representation for downstream tasks by addressing the human-robot domain discrepancy in pre-training.
5 Conclusions and Limitations
Learning generalizable representations from large-scale human data for robotic manipulation shows great potential. However, human-robot domain discrepancy remains a challenge that existing pre-training paradigms cannot effectively address. This work makes a preliminary attempt to tackle this challenge by introducing a new adaptation paradigm, which leverages paired human-robot data and proposes an efficient semantic alignment method. Through this approach, human-data pre-trained models can be explicitly and efficiently adapted to the robot domain without the need to be tailored for each downstream robotic environment. Experiments on 20 simulated manipulation tasks and five real-world manipulation tasks with different pre-trained models, demonstrate the effectiveness of our proposed method.
To motivate further research, we highlight several limitations of this work. First, exploring ways to measure domain discrepancies between human-data pre-training and downstream robotic tasks could guide the adaptation process and potentially improve outcomes. Second, this work focuses on adapting pre-trained models without revisiting the pre-training stage. Integrating paired human-robot data with out-of-domain human data for pre-training could lead to more advanced models and adaptation methods. Finally, further research is needed to understand how the scale and diversity of both pre-training data and paired human-robot data impact domain discrepancies in pre-training for robotic manipulation.
Acknowledgement
This work was supported by the National Natural Science Foundation of China (No. 62306257), and the Guangzhou Municipal Science and Technology Project (No. 2024A03J0619 & 2024A04J4390). This work was also supported by the Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2023A03J0008) and Education Bureau of Guangzhou Municipality.
Appendix
Appendix S1 Downstream Policy Learning on RLBench
Existing visual pre-training works often assess the learned representation in downstream environments using a single-task setting. However, our work diverges from this standard by also evaluating the adapted pre-trained models on the large-scale RLBench benchmark [32], where a single language-conditioned policy is learned to complete various tasks. Figure S1 shows examples of the 18 tasks and the corresponding human instructions.
On RLBench, existing works [10, 19, 24, 31, 60] usually develop sophisticated models to model the semantics of the robot’s multi-view observations and their correlations with the language commands. For example, RVT [19] utilizes four attention layers as the visual encoder to model intra-image relations and four more attention layers to model image-language correlations. In this work, we adopt the same design as RVT, but replace the visual encoder (i.e., four intra-image attention layers) with either an existing pre-trained model or our adapted one. Additionally, we employ just one attention layer to fuse the extracted image features and language features. Since RVT predicts the end-effector’s actions based on the features without down-sampling the spatial dimension, we discard all spatial down-sampling operations (e.g., max-pooling) in both the pre-trained models and our adapted models. Please note that to validate the effectiveness of our adaptation method, we freeze the visual representation of the pre-trained model or our adapted model while learning the downstream policy on RLBench.
Figure S1: Examples of the 18 RLBench tasks (front view) with corresponding human instructions (sourced from [43]).
Figure S2: Examples of the five real-world tasks are shown, with each row presenting an instance of the corresponding task. For each demonstration, we provide visual observations at six different timestamps.
Appendix S2 Real-world Experiments
Setups. For the real-world manipulation experiments, we use a 7-DoF xArm robot arm equipped with an Inspire gripper. Visual observations are captured by an Orbbec Femto Bolt (RGB-D) camera positioned in front of and to the upper right of the robot arm. The positions of the robot arm, working area, and camera remain fixed during data collection and policy testing. Additionally, we use a DJI Osmo Action 4 camera to record videos of the policy testing.
Data Collection. We design five different real-world tasks, namely, put fruit in plate, stack cups, put tennis in mug, hang mug, and put item in box. For each task, we collect 40 human teleoperation demonstrations for training. Figure S2 demonstrates some examples of the collected tasks. For each demonstration, we manually move the robot arm and change the states of the gripper (i.e., open or close) to complete the target task. We record these operations and replay them to record the demonstration. We simultaneously record the robot arm end-effector state (i.e., positions in the x-axis, y-axis, and z-axis, and rotations in roll, pitch, and yaw), gripper state, and Orbbec camera RGB stream with an image size of 1280$\times$960.
Model Designs. We train our manipulation policy network under a single-task setting, utilizing the ACT [76] framework for policy learning***We follow the implementation of https://github.com/Shaka-Labs/ACT.. In addition, following RVT [19] which predicts the next key-action, we predict the following key-actions of the end-effector. The visual backbone of the network is replaced with either the pre-trained models or our adapted models. During both training and testing, the RGB images are resized to $320\times 240$.
Models learned params. pen relocate Averaged
R3M 0M (frozen:25M) 78.0 70.0 74.0
R3M-Align-L 1.6M 81.3 81.3 81.3 (+7.3)
@@ Acknowledgement
Table S1: Success rates of two tasks in Adroit. Removing the language-guided feature enhancement will degrade the model’s performance.
Appendix S3 Additional Ablations
In this work, our adaptation method uses task description features as queries to better capture task-aware semantics from video features. As shown at the bottom of Table S1, by removing this operation, the adapted model, i.e., R3M-Align-L w/o lang., will result in performance degradation. This demonstrates that the language-guided feature enhancement is effective in promoting human-robot semantic alignment.
[…]
R3M-PreT R3M-ClS R3M-Align (Ours)
78.1 77.5 81.3
Table S2: Comparisons between models when training with full data.
R3M-PreTA R3M-ClSA R3M-Align (Ours)
77.2 76.9 81.3
Table S3: Comparisons between models when training with Adapters.